OpenStudy (anonymous):

Look at the numbers below. 8, 7, 5, 4 Which pair of numbers has an odd product? A. 5 and 4B. 7 and 4C. 7 and 5D. 8 and 7

OpenStudy (anonymous):

7 & 5

OpenStudy (anonymous):

5*4=20, not odd 7*4=28, not odd 7*5=35 odd, 8*7=56, not odd. C. 7 and 5